OpenStudy (anonymous):

Send it quick mail house mailed five packages with a mean weight of 6.7 LBS. Suppose the mean weight of four of these packages is 7.2 LBS. What is the weight in LBS of the fifth package? A.) 3.35 B.) 4.7 C.) 6.95 D.) 8.7

OpenStudy (anonymous):

total weight of four packages = 7.2 * 4 = 28.8 mean for five packages: 6.7= (total weight of four packages + weight of 5th package) / 5

OpenStudy (anonymous):

5.76?

OpenStudy (anonymous):

6.7= (total weight of four packages + weight of 5th package) / 5 33.5=28.8+ weight of 5th package